OpenStudy (anonymous):

Express answer in exact form. Find the area of one segment formed by a square with sides of 6" inscribed in a circle. (Hint: use the ratio of 1:1:√2 to find the radius of the circle.)

OpenStudy (anonymous):

radius of circle = sqrt(2)*3 area of circle = pi*(sqrt(2)*3)^2 area of segment = (area of circle - area of square)/4 area of segment ~ 5.137 inches^2